OpenStudy (anonymous):

5/ 2 square root 2-1

OpenStudy (anonymous):

2 minus one?

OpenStudy (anonymous):

it is the square root 2... then minus 1 im not sure how to put it on the cal..

OpenStudy (anonymous):

oh oops.

OpenStudy (anonymous):

*calcula

OpenStudy (anonymous):

wait is it this:\[\sqrt{2-1}\] or this: \[\sqrt{2}-1\]?

OpenStudy (anonymous):

yes it will be the 2nd one but with a 2 in front of it

OpenStudy (carson889):

The latter. And putting it into the calc, it should look like this: ((5/2)*((2)^(1/2)))-1

OpenStudy (carson889):

\[(\frac{ 5 }{ 2 }\sqrt{2}) -1\] = apprx 2.5355

OpenStudy (anonymous):

2.7346 is the answer

OpenStudy (anonymous):

\[\frac{ 5 }{ 2\sqrt{2}-1 }\] is it this one?

OpenStudy (anonymous):

yes! hehe

OpenStudy (anonymous):

so you got the answer?